Calculate the interquartile range of the given data.
53;4;14;55;2;34;27;48;51;34;37;8;4;35;34
Calculate the interquartile range of the given data. 53, 4, 14, 55, 2,34, 27,48, 51,34,37,8,4.35,34
Answer #1

Solution: -

We have given data as follows,

53 04 14 55 02 24 27 48 51 34 37 08 04 35 34

Insert the given data in excell sheet,

Calculations: -

Here we have to find Interquartile range for given data.

i. e. Q3-Q1=(3rd quartile) - (1st quartile)

But

3rd quartile is always 75 th percentile, and

1st quartile is always 25 th percentile.

Hence,

For finding Q3 and Q1 by giving command in MS-excell

As

=PERCENTILE (array, k) ### where, array: select all data set.

k: percentile value, k=0.25 for Q1

k=0.75 for Q3

Therefore,

Q3=42.5 and Q​​​​​​1=11

Q​​​​​3-Q1=42.5-11 = 31.5

Results: -

Interquartile range for given data set is  31. 5